OpenStudy (jack1):

how would i find the equation of a line that passes through the following points please:

OpenStudy (jack1):

2, 0.6935 3, 1.0986 6, 1.7918

OpenStudy (jack1):

i know it's a quadratic, i just don't know how to find a, b and c from it

OpenStudy (ddcamp):

Solve the system of 3 equations: 0.6935 = A(2²) + B(2) + C 1.0986 = A(3²) + B(3) + C 1.7918 = A(6²) + B(6) + C

OpenStudy (jack1):

genius, cheers for that